A transformer supplies 31kVA to a 25kW AC motor. Assume the motor is 100% efficient. What is the power factor?
Answer #1

Motor Efficency is 100%

So, Input Power is totally transfered to the Load without any losses

Input Power = Output Power

Input Power = KVA Rating * Power factor

KVA Rating given = 31 KVA

Output Power given = 25 KW

25 = 31*(Power factor)

Power factor = 25/31 = 0.8064